OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by ten.

OpenStudy (anonymous):

\[ \frac{ \pi }{ 10 }\]

OpenStudy (anonymous):

I know that cot is \[\frac{ adjacent }{ opposite }\]